OpenStudy (anonymous):

if Q(x) = x2-5x+1 find Q(2+h) - Q(2) /h

OpenStudy (anonymous):

didn't we just do this or am i mistaken?

OpenStudy (anonymous):

i didn't understand the first one i changed 1 to 2

OpenStudy (anonymous):

Sorry i'm confused :|

OpenStudy (anonymous):

ok this time we do it with numbers what is \(Q(2)\) ? we need that number

OpenStudy (anonymous):

the Q(x) multiplied by 2?

OpenStudy (anonymous):

no the function evaluated at \(x=2\)

OpenStudy (anonymous):

\(Q(x)\) means "Q of \(x\)" not \(Q\times x\) similarly \(Q(2)\) means "Q of 2" replace \(x\) by \(2\) \[Q(x)=x^2-5x+1\] \[Q(2)=2^2-5\times 2+1\]

OpenStudy (anonymous):

ohhh i seee Thank You :D

OpenStudy (anonymous):

you also need to find \(Q(2+h)\) which will have an \(h\) in it, whereas \(Q(2)\) is a number

OpenStudy (anonymous):

did you find \(Q(2)\) ?

OpenStudy (anonymous):

im having a hard time.. this is my first time :|

OpenStudy (anonymous):

did you compute this \[Q(2)=2^2-5\times 2+1\]?